# Optimas: Optimizing Compound AI Systems with Globally Aligned Local Rewards

### v2

@@ 4.3 Optimization with Globally aligned Local Reward Functions
Given a globally aligned LRF $r_{k}$, we perform local optimization on each component $C_{k}$ to improve its configuration $\mathbf{v}_{k}$. Specifically, we solve:
In practice, explicitly setting the trust region threshold $\delta$ can be difficult due to heterogeneous configuration types (e.g., continuous weights vs. discrete tokens). Instead, we adopt a conservative number of update steps to restrict the magnitude of change during each iteration.
Local Optimization. Since each LRF is associated with a specific component, Optimas can flexibly apply a specialized optimization method for each component. Further implementation details are provided in Appendix E. Concisely,
  • For textual prompts, we use prompt optimization algorithms such as OPRO [17] that ranks candidate prompts by their average local reward and select the best-performing one.
  • For components that are trainable models (e.g., an LLM), we apply reinforcement learning—such as Proximal Policy Optimization (PPO) [22]—using the LRF as the critic.
@@
Summary (Figure 3). Optimas leverages offline training to learn a set of LRFs that are well-aligned with the global reward under the initial system configuration. During optimization, Optimas iteratively conducts local optimization to update individual component configurations with high data efficiency, and adapts the corresponding LRFs to align with the evolving system using minimal data.
@@ 4.4 Theoretical Insights
We prove that the local–global alignment property (4) holds for the LRFs constructed in Section 4.1.
@@
Under regulartiy conditions, the maximizer of (5) satisfies the local-global alignment property (4). In addition, maximizing $r_{k}\bigl{(}x_{k},\,C_{k}(x_{k};\mathbf{v}_{k})\bigr{)}$ over $\mathbf{v}_{k}$ and maximizing $R(x,f(x;\mathbf{v}_{-k})\mid C_{k}(x_{k};\mathbf{v}_{k}))$ over $\mathbf{v}_{k}$ will yield the same solution.
We defer the formal statement for this theorem to appendix A. Since solving (2) is generally challenging, below we introduce some regularity conditions to make the convergence analysis tractable.
As the configurations $\mathbf{v}$ are heterogeneous, where some of the coordinates are discrete, and some are continuous, without loss of generality, we assume the first $M$ configurations $\mathbf{v}^{(1)}=\{\mathbf{v}_{1},...,\mathbf{v}_{M}\}$ are continuous, and the last $(K-M)$ configurations $\mathbf{v}^{(2)}=\{\mathbf{v}_{M+1},...,\mathbf{v}_{K}\}$ are discrete. We write the objective function $\mathbb{E}_{x\sim\mathcal{D}}\bigl{[}R(x,f(x;\,\mathbf{v})\bigr{]}$ as $l(\mathbf{v})=l(\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{K}):=l(\mathbf{v}^{(1)},\mathbf{v}^{(2)})$.
@@ 5 Experiments
We evaluate Optimas against five baselines on five real-world compound AI systems across diverse domains, aiming to validate its ability to improve system performance. Below, we summarize the dataset settings, baselines, and evaluation metrics, with further details provided in the appendix.
Benchmarks & Evaluation (Appendix B). We briefly introduce the benchmarks and task metrics.
@@
  • Amazon [51]: A behavior-driven recommendation task based on Amazon product interactions, evaluated using accuracy to measure if the predicted next item matches the ground-truth item.
@@
  • PubMedQA [52]: A clinical classification dataset derived from PubMed abstracts [53], evaluated by accuracy, defined as the proportion of predictions that exactly match the ground-truth labels.
@@
  • STaRK-Prime [54]: A retrieval benchmark over semi-structured biomedical corpora, evaluated using Mean Reciprocal Rank (MRR), which averages the inverse rank of the first correct answer.
@@
  • HotpotQA [55]: A multi-hop question answering dataset, evaluated using the F1 score between predicted and ground-truth answers.
@@
  • BigCodeBench [56]: The instruction split of BigCodeBench for self-verifying code generation, evaluated using pass rate: the percentage of final generated programs that pass the unit tests.
@@
Compound Systems (Figure 2, cf. Appendix C). We design a compound system per benchmark with diverse and common patterns for agentic systems [^1] https://www.anthropic.com/engineering/building-effective-agents. All systems are accessible in our code repository.
Baselines (Appendix D). We compare Optimas against five comprehensive baselines:
  • LLMSelector [8]: A lightweight policy selects the best LLM per component via model routing, without updating prompts or other variables. This method is only applicable on PubMedQA.
  • Hierarchical Behavior Cloning (HBC) [57]: A hierarchical imitation learning approach that optimizes each component to produce outputs similar to those that lead to high global rewards.
  • TextGrad [1]: A gradient-based prompt tuning method using estimated gradients from black-box LLMs to improve prompt efficacy.
  • DSPy [2, 23]: A prompt optimization framework using the MIPRO algorithm that jointly refines module-level instructions and few-shot demonstrations.
(a) Amazon (Fig. 7(a))
(b) STaRK-Prime (Fig. 7(b))
(c) HotpotQA (Fig. 7(c))
@@
Figure 4: Global reward of the three compound AI systems over local optimization iterations. For conciseness, we show only the steps of local optimization that lead to an increase in global reward on the validation sets. The annotations indicate the components optimized at each step.
@@
Figure 5: Case Study on HotpotQA: (a) Before local optimization, we collect outputs from Answer Generator, where we compute each output’s local reward from LRF and global reward, demonstrating that the local reward highly correlates with the global reward. (b) Optimas optimizes prompts on Answer Generator to maximize local rewards, achieving higher global reward. (c) Interpretation: LRF learns to prefer concise answers from the local-global alignment.
5.1 Final Global Reward and Optimization Dynamics
Consistent Gains vs. Strong Baselines (Table 2). We compare Optimas with the baselines under similar numbers of system runs. Under this controlled data cost (Table 3), Optimas consistently improves global rewards across all five compound systems, achieving an average relative improvement of 11.92% compared to the best baseline.
While the strong baseline DSPy shows notable improvements on some datasets—for instance, a 21.6% gain on HotpotQA—their performance may be inconsistent and can even degrade the system-a 14.3% drop on the Amazon dataset. For LLMSelector, the total system runs cost is 2.8k which is 3x more expensive than Optimas. The comparison highlights the advantage of using globally aligned reward functions to guide local optimization.
Progressive gains during Optimas optimization (Figure 4). Further, we visualize the changes in global reward across optimization iterations. Within a small number of iterations, Optimas achieves a substantial average improvement of 41.7% over the initial global reward on the validation sets, using lightweight and data-efficient updates. Specifically, by updating each component during local optimization, Optimas approximates the optimal system configuration. Full configuration traces are provided in Appendix E.
5.2 In-depth Study of Optimas
We analyze Optimas to provide insights and understand how its core mechanism—local-global alignment—contributes to improvements in overall system global reward.
@@ 5.2 In-depth Study of Optimas
(1) The quality of local-global alignment (Table 4). To assess alignment quality, we compute pairwise ranking accuracy: the probability that an output with higher global reward receives a higher local reward than an output with lower global reward. This metric reflects how well the learned LRFs aligns with global rewards. We compare against a baseline, LLM Judge, which prompts a gpt-4o model to rank outputs based on 20 in-context input-output examples—10 associated with high global rewards and 10 with low global rewards. This approach resembles prior methods such as TextGrad, which rely on few-shot reasoning over textual patterns.
In Table 4, LLM Judge performs closer to random guessing, likely due to the diversity and stochasticity of intermediate outputs in compound systems. These factors make it difficult to infer output quality based solely on static textual patterns. In contrast, the learned LRFs achieve a substantially higher average pairwise accuracy of 77.96%, indicates the quality of the local reward produced by our LRFs compared to textual reasoning. Moreover, LRFs are trained to internalize the local-global alignment within their weights, eliminating the reliance on limited in-context examples or textual patterns, and enabling more precise alignment with global reward.
(3) Insights learned by LRFs during local-global alignment (Figure 5). We examine the LRF learned for the Answer Generator in the HotpotQA system. Figure 5(a) shows a positive correlation between the local reward and global reward of each predicted answer, indicating that our LRF captures how each local output contributes to the global reward.
Figure 5(b) illustrates how local optimization with LRF refines the prompt. Initially, the prompt instructs the generator to “directly answer the query with a short answer,” yielding a local reward of 0.17 and a global reward of 0.32. After optimization with LRF, the instruction becomes more specific: “concise answer to the query in no more than 10 words …,” increasing the local reward to 0.22 and the global reward to 0.36. This improvement can be attributed to the clear negative correlation between answer length and local reward, as shown in Figure 5(c), where shorter responses tend to be rewarded more. Since correct answers in HotpotQA are often concise, this correlation naturally guides the system to enforce brevity through prompt optimization.
(3) Correlation between alignment quality and global reward (Figure 6 (a) and (b). To better understand how alignment quality affects global reward, we conduct controlled experiments on the Amazon and HotpotQA systems. For each system, we select a key component (Candidate Profiler for Amazon and Answer Generator for HotpotQA) and train local reward models with varying alignment quality, measured by pairwise ranking accuracy. Pairwise ranking accuracy quantifies how often an output with higher global reward is assigned a higher local reward compared to an output with lower global reward.
We then perform single-step local optimization using each trained model and measure the average global reward achieved after updating only that component. Figure 6 shows a strong positive correlation between the alignment quality of the local reward model and the corresponding improvement in global reward. Specifically, local reward models with higher alignment quality consistently produce larger global reward gains, underscoring that accurate local-to-global alignment is critical for effective and reliable local optimization. Conversely, lower-aligned local rewards may misguide optimization and lead to negligible or even negative improvements in global reward.
(a) Amazon alignment
(b) HotpotQA alignment
(c) Data sensitivity study
Figure 6: Part 1: Correlation between alignment quality and global reward on (a) Amazon and (b) HotpotQA. Local reward models with varying alignment quality—measured by pairwise ranking accuracy—are used to optimize selected components; a strong positive correlation is observed, with higher alignment consistently yielding larger global reward gains. Part 2: Impact of training-data size on evaluation accuracy for HotpotQA (c). Optimas remains robust even with limited data, showing only a 5.19% absolute drop when trained on just 12.5% of the data compared with the full dataset, demonstrating efficient data usage.
(4) Ablation on Data Size (Figure 6 (6(c))). To evaluate the data efficiency of Optimas, we examine how the amount of training data (system runs) affects the performance of learned LRFs. We train local reward models on the HotpotQA system using varying percentages of available training data: 12.5%, 25%, 50%, 75%, and 100%. Each model is trained for 3 epochs, and performance is measured by pairwise ranking accuracy—how often an output with higher global reward is assigned a higher local reward compared to an output with lower global reward.
Optimas maintains reasonable performance across different data sizes. Even with only 12.5% of the data, the system achieves 65.46% accuracy (92.7% of full-data performance), while 25% of the data yields 67.22% accuracy (95.1% of full-data performance). The diminishing returns beyond 25% data suggest that the core patterns for local-global alignment can be captured efficiently from limited preference data. This data efficiency is particularly valuable in practice, where collecting extensive preference data through system runs can be computationally expensive. Combined with the reduced computational cost of local optimization, these results demonstrate that Optimas provides a practical solution for optimizing complex compound AI systems even in resource-constrained environments.
@@ 6 Conclusion
We presented Optimas, a unified framework to optimize compound AI systems end-to-end. Its key insight is to maintain one Local Reward Function per component, aligned with the system’s global performance. This design allows every component, whether a fine-tunable LLM, API-based LLM, tool hyperparameters, or model selection, to be optimized locally while still improving the overall system. Through extensive experiments on five real-world tasks, we demonstrated that globally aligned local optimization outperforms strong baselines, including specialized prompt-tuning, hierarchical behavior cloning, and model selection approaches. While our approach effectively reduces system-wide runs, a current limitation is that we choose which component to optimize using a simple gap-based heuristic; developing a more principled or learned scheduler for selecting both the component is an important direction for future work. We believe Optimas will serve as a general, data-efficient approach for continually refining multi-component AI systems in diverse application domains.
Appendix A Theoretical Analysis

### v3

@@ 4.3 Optimization with Globally aligned Local Reward Functions
Local Optimization. As each component has its own LRF, Optimas flexibly applies a specialized optimization method for each component. See details in Appendix F. Concisely,
  • For textual prompts, we use prompt optimization algorithms such as OPRO (Yang et al., 2024) that ranks candidate prompts by their average local reward and select the best-performing one.
  • For components that are trainable models (e.g., an LLM), we apply reinforcement learning—such as Proximal Policy Optimization (PPO) (Schulman et al., 2017)—using the LRF as the critic.
@@
Overall algorithm (Figure 3). Start with the initial system configuration, Optimas leverages initial reward modeling to learn a set of LRFs that are well-aligned with the global reward. At each iteration optimization, Optimas randomly selects a component to optimize, adapts its LRF to align with the evolving system using minimal amount of data, and conducts local optimization to obtain a new configuration. To prevent potential cascading errors, the new configuration is accepted only if it improves the system performance on a small validation set. Since the optimization with LRFs is conducted locally, the number of system runs to achieve a high global reward is reduced, as we later show in the experiments. The detailed algorithm is provided in Appendix A.
@@ 4.4 Theoretical Insights
We prove that the local–global alignment property holds for the LRFs constructed in Section 4.1.
@@
Under regulartiy conditions, the maximizer of equation 4 satisfies the local-global alignment property (equation 3). In addition, maximizing $r_{k}\bigl(x_{k},\,C_{k}(x_{k};\mathbf{v}_{k})\bigr)$ over $\mathbf{v}_{k}$ and maximizing $R(x,f(x;\mathbf{v}_{-k})\mid C_{k}(x_{k};\mathbf{v}_{k}))$ over $\mathbf{v}_{k}$ will yield the same solution.
We defer the formal statement for this theorem to appendix B. Since solving equation 1 is generally challenging, we introduce some regularity conditions to make the convergence analysis tractable.
As the configurations $\mathbf{v}$ are heterogeneous, where some of the coordinates are discrete, and some are continuous, without loss of generality, we assume the first $M$ configurations $\mathbf{v}^{(1)}=\{\mathbf{v}_{1},...,\mathbf{v}_{M}\}$ are continuous, and the last (KM)(K-M) configurations $\mathbf{v}^{(2)}=\{\mathbf{v}_{M+1},...,\mathbf{v}_{K}\}$ are discrete. We write the objective function $\mathbb{E}_{x\sim\mathcal{D}}\bigl[R(x,f(x;\,\mathbf{v})\bigr]$ as $l(\mathbf{v})=l(\mathbf{v}_{1},\mathbf{v}_{2},...,\mathbf{v}_{K}):=l(\mathbf{v}^{(1)},\mathbf{v}^{(2)})$.
@@ 5 Experiments
We firstly summarize the datasets, baselines, and metrics, with details provided in the appendix.
Benchmarks & Evaluation (Appendix C). We evaluate Optimas on five real-world tasks:
@@
  • Amazon (Jin et al., 2024): A behavior-driven recommendation task based on Amazon products, evaluated using accuracy to measure if the predicted next item matches the ground-truth item.
@@
  • PubMedQA (Jin et al., 2019): A clinical classification dataset derived from PubMed abstracts (National Center for Biotechnology Information (2024), NCBI), evaluated by accuracy, defined as the proportion of predictions that exactly match the ground-truth labels.
@@
  • STaRK-Prime (Wu et al., 2024b): A retrieval benchmark over semi-structured biomedical corpora, evaluated using Mean Reciprocal Rank (MRR).
@@
  • HotpotQA (Yang et al., 2018): A multi-hop question answering dataset, evaluated using the F1 score between predicted and ground-truth answers.
@@
  • BigCodeBench (Zhuo et al., 2024): The instruction split of BigCodeBench for self-verifying code generation, evaluated using pass rate.
@@
Compound Systems (Figure 2, cf. Appendix D). We design a compound system per benchmark with diverse and common patterns for agentic systems. All systems are accessible in our code repository.
Baselines (See Appendix E for details). We compare Optimas against five baselines: Unoptimized, LLMSelector (Chen et al., 2025b), Hierarchical Behavior Cloning (HBC) (Le et al., 2018), TextGrad (Yuksekgonul et al., 2025), DSPy (Khattab et al., 2023; Opsahl-Ong et al., 2024).
(a) Amazon system
(b) STaRK-Prime system
(c) HotpotQA system
@@
Figure 4: Global reward and configuration updates of the three compound AI systems over the optimization iterations. For conciseness, we only show the local optimization steps that lead to an increase in global reward on the validation sets. The annotations show the optimized components.
5.1 Performance of the systems during and after optimization.
Takeaway 1: Optimas leads to consistent and substantial improvements (Table 2). We compare Optimas with the baselines under similar numbers of system runs. Under this controlled data cost (Table 3), Optimas consistently improves global rewards across all compound systems, achieving an average relative improvement of 11.92% compared to the best baseline.
While the strong baseline DSPy shows notable improvements on some datasets (e.g., a 21.6% gain on HotpotQA), their performance may be inconsistent and can even degrade the system (e.g., a 14.3% drop on the Amazon dataset). For LLMSelector, it requires $2.8k$ times of forwarding through the entire system, which is $3x$ more expensive than Optimas.
Takeaway 2: Local optimization improves global rewards (Figure 4). We study how the configurations change in the local optimization. Within a small number of iterations, Optimas achieves a substantial average improvement of 41.7% over the initial global reward on the validation sets, using lightweight and data-efficient updates. Interestingly, we observe a mixed updates on prompts, model parameters, and hyperparameters, which can lead to improved global reward. For example, updating the prompt of Text Scorer in the 9-th iteration improves global reward from $0.49$ to $0.56$. Among these cases, they involve optimizing different components to achieve the highest global reward empirically, showing the importance of being able to optimize different types of configurations.
5.2 Why and How Optimas Works: Alignment, Interpretability, and Efficiency
@@
We conduct extensive in-depth study to understand the mechanism in Optimas framework.
@@
Takeaway 3: Optimas yields high local-global alignment quality (Table 4). To measure alignment quality of LRFs, we compute pairwise ranking accuracy: the probability that an output with higher global reward receives a higher score than an output with lower global reward. This reflects how well the learned LRFs aligns with global rewards. We compare against a LLM Judge, which prompts a gpt-4o model to score the outputs of components based on 20 in-context examples. This approach is similar to prior methods such as TextGrad, which rely on few-shot reasoning over textual patterns. In Table 4, LLM Judge performs closer to random guessing, due to the diversity and stochasticity of components’s outputs that make it difficult to reason reliably. In contrast, our LRFs achieve substantially higher performances. Moreover, LRFs internalize the local-global alignment within their weights without relying on limited in-context examples, enabling more precise alignment.
Takeaway 4: Higher alignment quality usually leads to higher global reward (Figure 5). To understand how alignment quality affects global reward, we conduct controlled experiments, where we select a key component, apply local reward models with varying alignment quality to optimize it, and measure the average global reward achieved after updating the component. In Figure 5, we show a strong positive correlation between the LRF’s alignment quality and the global reward improvement.
(a) Candidate Profiler on Amazon
(b) Answer Generator on HotpotQA
Figure 5: Local reward models with varying alignment quality are used to optimize a selected component in each task, where we observe that higher alignment quality yields higher global rewards.
Figure 6: An interpretability study on what is learned by LRFs. With (a) a well-aligned LRF, we find that (b) the optimized prompt explicitly constrains the output length of the component. We attribute this to that (c) the LRF prefers short outputs, which is consistent with the use of F1 as global metric.
Takeaway 5: LRFs learn interpretable directions to improve global reward (Figure 6). An important aspect of system optimization is interpretability, i.e., if the configuration updates are reliable and understandable. We provide a study in Figure 6, where the LRF learns to favor concise answers. In fact, it is more feasible to interpret configuration updates with LRFs. Specifically, one can perturb the component outputs in certain ways, and observe the changes in local rewards to obtain insights.
Takeaway 6: Optimas is data and computational efficient. (Appendix G). We also conduct extensive studies to show the data efficiency with varying dataset size and computational efficiency with lightweight models. Due to space limitation, please refer to Appendix G.
@@ 6 Conclusion
Optimas is a unified framework to optimize compound AI systems with heterogeneous configurations. Optimas’ way to maintain globally aligned local reward functions allows every component, whether a fine-tunable LLM, LLM API, tools, or model selector, to be optimized locally while improving the overall system. On five real-world tasks, Optimas outperforms strong baselines, effectively optimizes components with different configurations, and exhibits high alignment quality and reliable interpretations. We believe Optimas will serve as a general, data-efficient approach for continually optimizing practical systems. For future work, we aim to further apply Optimas on even larger systems, with the goal of understanding complex reward modeling and scalability.
Ethics statement.
The applications of compound AI systems today include, but are not limited to, decision processes in socially and economically important domains, making their optimization an increasingly significant research problem. While our work aims to make optimization of such systems more data-efficient, robust, and interpretable, we recognize potential ethical concerns. In high-stakes settings such as healthcare or recommendation platforms, optimization could inadvertently amplify biases or propagate unsafe behaviors from individual components; our framework mitigates these risks by aligning local rewards with global performance, thereby reducing the likelihood of harmful emergent behaviors and creating opportunities to incorporate fairness- and safety-sensitive objectives. Another concern is malicious use of optimized compound systems; however, our approach relies on safety-aligned LLMs and tools, and we observe that aligning local and global objectives does not reduce the effectiveness of existing safeguards, while the decentralized optimization structure may even increase opportunities to detect misuse. Regarding data, our experiments rely solely on publicly available datasets or synthetic user interactions without any personally identifiable information (PII), and no private or sensitive user data were used. We believe this work contributes to the safe and responsible advancement of AI by providing both theoretical and practical insights into optimizing compound AI systems, and to promote further research in this direction we release all code, models, and benchmarks described in this paper.
Reproducibility statement
We provide full open access to our implementation code, including the original datasets, compound system implementations, and the Optimas framework at https://anonymous.4open.science/r/optimas-1308/. To support reproducibility, Appendix C (Dataset Details), Appendix D (System Details), Appendix E (Baseline Details), and finally, Appendix F (Optimas Details) further documents the datasets used and model architectures etc.. Together with our descriptions in the experiment section, we ensure that our data and code are fully accessible and our results are transparent and easily extensible.
Appendix A Algorithm
Algorithm 1 Component-wise optimization with reward-model adaptation
1:Input
[…]
5:Global reward $R(\cdot)$, preference dataset size $k$, total iterations $T$
6:Training dataset $\mathcal{D}$ and validation dataset $D_{v}$
[…]
9:Optimization loop
10:for $t=0,\ldots,T-1$ do
11:Scheduler: choose a component to optimize
12:  $i_{t}\sim\mathrm{Uniform}\big(\{1,\ldots,K\}\big)$
13:Reward-model adaptation
14:  $D_{t}\leftarrow\textsc{CollectPreferenceData}(\mathcal{D},\Theta^{t},i_{t},k)$ $\triangleright$ create $(x^{+},x^{-})$ pairs for $C_{i_{t}}$ based on $R(\cdot)$
15:  $\theta^{t+1}_{i_{t}}\leftarrow\textsc{RewardModelTrain}(\theta^{t}_{i_{t}},D_{t})$
16:  $\Theta^{t+1}\leftarrow\Theta^{t}$ with $\theta_{i_{t}}$ replaced by $\theta^{t+1}_{i_{t}}$
17:Local optimization for the chosen component $C_{i_{t}}$
18:  if $C_{i_{t}}$ is an LLM with prompts then
19:   $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}\leftarrow\textsc{PromptOptimization}(\mathbf{v}^{t}_{i_{t}},\theta^{t+1}_{i_{t}})$
20:  else if $C_{i_{t}}$ has trainable weights then
21:   $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}\leftarrow\textsc{PPOTrain}(\mathbf{v}^{t}_{i_{t}},\theta^{t+1}_{i_{t}})$
22:  else if $C_{i_{t}}$ has a hyperparameter configuration then
23:   $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}\leftarrow\textsc{HyperparameterSearch}(\mathbf{v}^{t}_{i_{t}},\theta^{t+1}_{i_{t}})$
24:  end if
25:  $\tilde{\mathbf{v}}\leftarrow\mathbf{v}^{t}$ with $\mathbf{v}^{t}_{i_{t}}$ replaced by $\tilde{\mathbf{v}}^{\,t+1}_{i_{t}}$
26:Validation
27:  if $\sum_{x_{v}\in D_{v}}R\!\big(x_{v},f(x_{v};\tilde{\mathbf{v}})\big)\;>\;\sum_{x_{v}\in D_{v}}R\!\big(x_{v},f(x_{v};\mathbf{v}^{t})\big)$ then
28:   $\mathbf{v}^{t+1}\leftarrow\tilde{\mathbf{v}}$;  $\mathbf{v}^{*}\leftarrow\mathbf{v}^{t+1}$
29:  else
30:   $\mathbf{v}^{t+1}\leftarrow\mathbf{v}^{t}$
31:  end if
32:end for
33:return $\mathbf{v}^{*}$
Appendix B Theoretical Analysis